### v1

# Demonstration of a reconfigurable ground to space quantum networking scheme

Abstract
We experimentally demonstrate a reconfigurable quantum network architecture suitable for integrating satellite links in metropolitan quantum networks. The network architecture is designed such that once a satellite is in range, it is configured in a multipoint-to-point topology where all ground nodes establish entanglement with the satellite receiver using time multiplexing to optimize long-distance transmission. Otherwise, the satellite up-link is rerouted to the ground nodes to form a pair-wise ground network. Leveraging both the time and frequency correlations of our photon-pair source, we demonstrate an increased coincidence-to-accidental ratio without additional resource overhead in a five-node network. To contextualize these experimental findings, we project their performance in a quantum key distribution scenario and outline a feasible route towards field deployment, using integrated photonics to enable network integration of up to 72 users.
keywords:
Quantum networks, satellite-based quantum information, quantum communication1 Introduction
The development of the quantum internet will enable advanced applications in quantum cryptography, quantum-enhanced metrology, and quantum computing [1, 2]. A fundamental building block of this vision is the ability to share quantum entanglement between distant nodes [3]. To-date, quantum networks have mostly been restricted to metropolitan distances and a small number of nodes with predominantly static implementations, constraining their scalability and adaptability for broader deployment [4, 5, 6, 7, 8, 9]. These limitations are largely due to photon absorption and scattering losses in optical fibres, which severely limit the range of entanglement distribution, as well as the significant infrastructure required to deploy and maintain terrestrial links over long distances. Satellites provide a promising platform to circumvent these challenges and distribute entanglement over global distances. However, there are significant challenges to integrate them seamlessly with terrestrial networks [10, 11, 12]. In particular, satellite links experience significant attenuation and intermittent connectivity during orbital passes [13, 14]. To overcome such limitations, some of the authors have proposed in [15] a reconfigurable network architecture. The network dynamically reconfigures its topology based on the satellite’s accessibility to optimize the limited availability of the satellite. During a satellite pass, the network operates in a multipoint-to-point configuration, enabling all ground nodes to establish entanglement with the satellite simultaneously. When the satellite is unavailable, the satellite uplink is rerouted via an optical switch to the ground nodes, transforming the network into a pair-wise connected ground topology. This reconfiguration ensures efficient and continuous operation across both configurations with minimal modifications to the infrastructure. Furthermore, the performance of each configuration is improved using time and frequency multiplexing.
In this paper, we experimentally demonstrate this network reconfiguration using correlated photon pairs generated from a periodically poled magnesium-oxide doped lithium niobate (PPLN) waveguide specifically designed for the Quantum Encryption and Science Satellite (QEYSSat) uplink scenario [16, 17]. In particular, non-degenerate broadband photon pairs at 1550 and 785 nm are produced from spontaneous parametric down-conversion to optimize both the ground-based optical fibre and the ground-to-space free-space link. Moreover, by leveraging time multiplexing, we demonstrate a linear performance increase for the up-link configuration with minimal satellite hardware requirements. In particular, a pulsed pump in conjunction with a frequency-to-time mapping assigns a distinct time delay to each spectral channel such that the signal can be demultiplexed from a single detector channel thus significantly reducing the satellite payload overhead [18, 3]. This resource alleviation extends to the ground configuration where only one detector per user is required. Indeed, the pulsed pump allows for channel discrimination by defining fixed time slots for the arrival of each channel at the detector thus reducing the number of accidental coincidences [4].
2 Experimental setup
The experimental setup is shown in Fig. 1. A mode-locked titanium-sapphire laser (Tsunami Spectra-Physics) with a repetition rate of 80 MHz produces 1042 nm light, which is converted in a BBO crystal (Castech Inc.) via second-harmonic generation and filtered down to a nm bandwidth centered around 521.4 nm. These downconverted photons then pump a 15 mm long PPLN waveguide (HC Photonics Corp.) such that broadband photon pairs at 787.5 nm (signal) and 1543.2 nm (idler) are produced via non-degenerate type-0 spontaneous parametric down-conversion (SPDC). PPLN waveguides offer high brightness, compactness, and compatibility with integrated devices [19]. After correction for losses, the brightness of the SPDC source was estimated to be pairs/sec/mW. A temperature controller was used to maintain the PPLN’s temperature at its phase matching value (30 ∘C) to maximize photon generation within the QEYSSat passband chosen for high atmospheric transmission and compatibility with high efficiency space-ready detectors [20]. Signal and idler photons are separated at a dichroic mirror and coupled into polarization-maintaining (PM) fibres. The signal fibre output is connected to two different detection scenarios depending on the satellite’s availability. For a satellite uplink, in either scenario the idler photons remain on the ground and are routed to the various users via wavelength division multiplexing (WDM) using concatenated telecom FBGs (O/E Land Inc.) and optical circulators (OZ Optics).
For the satellite configuration shown in Fig. 1a, a frequency-to-time mapping (FTM) unit enables a time-resolved detection of the distinct frequency channels on the same Si-APD detector. This FTM is implemented using a fibre Bragg grating array that inscribes a distinct time of arrival to each spectral channel, enabling multipoint-to-point communication with the satellite as depicted in Fig. 1c. In contrast, for the ground network, in Fig. 1b, a fully-connected topology between Alice, Bob and Charlie is established by wavelength division multiplexing (WDM) using individual FBGs. To optimize detection efficiencies around both and nm, Bob and Charlie were equipped with a Si-APD and SNSPD detector (Quantum Opus). This requirement could be overcome by using a broadband detector [21], thus eliminating the need for multiple detectors tailored to specific wavelengths, thereby reducing the system complexity. To demonstrate this scalable approach, the and nm photons, corresponding respectively to channels and in Fig. 1d, are combined into the same single-mode fibre (Thorlabs SMF-28) using a 1550/780 nm wavelength division multiplexer (Seagnol Photonics) and jointly recorded on Alice’s SNSPD detector. Although the 780 nm photons will be slightly multimode and suffer significant optical loss in the telecom fibre, Meyers-Scott et al. [22] demonstrated the high-fidelity transmission is nonetheless possible and Vinet et al. [15] showed that positive nonzero key rates can be generated for up to 32 km, enough to span a metropolitan area network.
3 Experimental results
To demonstrate the multiplexing advantage, we measured the coincidence-to-accidentals ratio (CAR) in both scenarios using up to four correlated channels specified within the broadband spectrum of the SPDC source, see Fig. 2. The CAR was calculated from , where and are the coincidence rates at matched and un-matched slots as defined in [23]. The former corresponding to true coincidences caused by photons generated in the same pump pulse whereas the latter represents accidental coincidences arising from multi-pair emissions or detector dark counts. Exploiting time and frequency multiplexing at the satellite and ground nodes, accidental coincidences, due to uncorrelated channel detections, are filtered leading to a CAR improvement in accordance with the theoretical expectation as shown in Fig. 3(a). More explicitly, one can calculate the CAR from the correlated photon pair rate (), the transmittance of the signal ( and idler channels () and the average signal and idler count rate according to [23]:
Assuming multiplexed channels, resulting in a linear CAR improvement [24]. To limit the cross-talk between the channels, 1 nm guard bands between the signal channels were used to limit their spectral overlap.
From the CAR, we approximate the two-photon interference visibility using the relation derived in [25]. The multiplexing scheme enhances the visibility in Fig. 3(b), allowing increased coincidence rates at higher pump powers. To contextualize these results for practical applications, we estimate the quantum bit error rate (QBER) according to and the asymptotic secure key rate (SKR) for the entanglement-based quantum key distribution (QKD) protocol outlined in [26, 27]. These results are respectively presented in Figs. 3(c)-3(d). Again, multiplexing allows for a reduction of the QBER and a higher attainable secure key rate. This method thus presents a scalable approach to enhance the entanglement generation rate in a multipoint-to-point quantum link during the satellite pass. Moreover, additional users can be incorporated to the ground network without compromising the quality of service or pairwise entanglement generation rate - and with minimal resource overhead. By allocating a distinct temporal window to each channel within the pulse train [4], each user can unambiguously discriminate between frequency channels using a single detector. Accordingly, coincidences, for the ground configuration at high pump power in Fig. 3(a), were measured with only one detector channel per user. Despite the limited detection efficiency of the SNSPD at 780 nm (), temporally resolving each channel still yields a twofold CAR improvement.
4 Discussion
Photon bandwidth and detector timing resolution underpin, respectively, the capacity for wavelength and time multiplexing. Consequently, for the satellite configuration, the number of channels is bounded by , where is the pulse repetition rate and is the timing resolution of the detection system, including synchronization-induced jitter. There is a clear compromise between the pulsed laser repetition rate and the number of available time-multiplexed channels. This trade-off is illustrated in Fig. 4 for our source, assuming a pump power of 100 W, channel loss of 30 dB, detector dead time of 1 s, 1000 counts per second dark count rate and timing jitter of 130 ps.
Owing to the requirement for spectral compatibility with the QEYSSat payload, the signal bandwidth of the ground configuration is restricted to the satellite acceptance window of 780–795 nm. Nevertheless, in contrast to the existing QEYSSat source, which is limited by its intrinsically narrow spectral bandwidth and supports essentially single-channel operation, the proposed configuration leverages the full 15 nm acceptance window. Assuming 100 GHz spacing between channels, this allows for up to 72 parallel channels, markedly enhancing the performance during the satellite pass [28]. The network could be further optimized by using flexible grid technology [29]: employing a wavelength selective switch or reconfigurable optical add-drop multiplexer [30] would not only simplify the experimental design but also enable adaptive bandwidth management beyond the two configurations showcased in Fig. 1. Moreover, under flex-grid channel allocation, entanglement distribution losses are fixed and independent of the number of users. This provides a significant scalability advantage over the standard WDM approach.
Compact, adaptable, and cost-effective technologies will be essential to scale quantum networks toward global deployment. Recent work on integrated photonic entangled photon sources has shown significant improvements in stability, energy efficiency, and miniaturization paving the way toward scalable quantum networking systems [31, 32, 33]. Motivated by these efforts, we propose in Fig. 5 an integrated entangled photon source design. The design emphasizes practicality and compatibility with satellite-enabled quantum networks. To minimize size and cost, the titanium-sapphire in Fig. 1 is replaced by a gain-switched laser diode that produces broadband optical pulses. The frequency-to-time mapping is applied on the pump field using a Y-branched digital optical switch [34] in combination with a spiral waveguide containing an array of on-chip Bragg gratings. By encoding the FTM on the pump photons, the insertion loss associated to its optical components can be compensated. After reflection, the temporally stretched pulse is directed through a thermally tuned unbalanced Mach-Zehnder interferometer, where a coherent superposition of early and late time-bins is encoded. The resulting pulse is then coupled into a PPLN waveguide, where time-bin entangled signal and idler photon pairs are generated via SPDC. These photon pairs are subsequently separated by a band-wavelength division multiplexer (BWDM). A second optical switch toggles between the satellite and ground configurations, rerouting the signal to a WDM for the latter. To evaluate the required guard bands for suppressing cross-talk in the frequency-to-time mapping stage, Fig. 6 shows a simulated joint spectral intensity calculated from the SPDC process of the proposed integrated entangled photon source. Using parameters consistent with the demonstration presented in section 2, a nm pump was divided into four frequency channels, each nm wide. nm spectral gaps were inserted between the Bragg grating reflection bands. This configuration was found to reduce interchannel cross-talk to below dB, ensuring that the down-converted photon pairs generated from different pump channels remain spectrally and temporally well separated.
This design could easily be extended to additional channels by increasing the number of Bragg gratings in the FTM unit. Such a source would enable quantum communication nodes that are portable, power-efficient, and suitable for deployment in real-world settings. The design consolidates the frequency-to-time mapping [35, 36], phase control, nonlinear photon-pair generation [37, 38, 39] and demultiplexing onto a single photonic chip [40, 41], significantly reducing size, alignment overhead, and sensitivity to environmental noise.
Declarations
Acknowledgements The authors acknowledge helpful discussions with Ramy Tannous, Julien Beaudoin Bertrand, and Mamoun Wahbeh.
Funding
This work is supported by the High Throughput and Secure Networks (HTSN) challenge by the National Research Council Canada, project HTSN-630. The authors further acknowledge support from the Natural Sciences and Engineering Research Council of Canada (NSERC), the Canadian Foundation for Innovation (CFI), the Ontario Research Fund (ORF). SV would like to thank the NSERC CGS-D for personal funding.
Conflict of interest/Competing interests
The authors declare no conflicts of interest related to this article.
Appendix A Materials and Methods
The total optical loss in the experimental setup arises from several distinct stages. At the chip interface, light was coupled into and out of the waveguide with 532 nm and 1550 nm polarization-maintaining lensed fibres (OZ Optics), incurring a coupling loss of approximately 4.5 dB. Following this, a free space bridge separated the signal and idler photons, with measured coupling efficiencies of 44 (3.57 dB loss) for idler photons and 16 (7.96 dB loss) for signal photons. The FTM unit introduced a significant insertion loss of 10 dB, due to the loss of 3 dB from the optical circulator and 7 dB attributed to the fibre Bragg grating (FBG) array. On the idler path, demultiplexing was performed using a cascaded assembly of circulators and FBGs, resulting in insertion losses of 4.86 dB, 7.73 dB, 10.48 dB, and 8.51 dB for the Alice, Bob, Charlie, and Dana channels, respectively. The 1550/780 WDM unit used for the ground configuration in Fig. 1b had an insertion loss of 0.97 and 3.02 dB at 1550 and 780 nm respectively.